Let the random variable Z follow a standard normal distribution. Find P(-2.35 < Z< -0.65). Your Answer:
Answer #1

(Date solution : given that, P(-2.35 L22-0.65) = P(22-0.65) - P(Z <-2.35) - using standard normal table. = 0.2598 - 0.0094 Pr